OpenStudy (bahrom7893):

Differential equations: Use the method of inspection to solve: y*dx + ( 2x^2y - x )*dy = 0

OpenStudy (anonymous):

method of inspection?

OpenStudy (bahrom7893):

yeah brb

OpenStudy (bahrom7893):

y*dx + x(2xy - 1)dy = 0 y + x(2xy-1)*(dy/dx)=0 y/[x(2xy-1)] + (dy/dx) = 0 dy/dx = - y / ( 2x^2y - x ) Does that look similar to any derivative equations?

OpenStudy (anonymous):

hmmm

OpenStudy (bahrom7893):

wait im lookin thru my stuff i think i can solve this afk..

OpenStudy (anonymous):

ok let me know

OpenStudy (bahrom7893):

http://i54.tinypic.com/4rtg0j.png So there are a couple of Derivative equations here, can that be simplified to any of them, this is basically algebra, but i don't see anything similar

OpenStudy (anonymous):

any other questions?

OpenStudy (anonymous):

ok ill figure it out one sec

OpenStudy (anonymous):

bahrom, why did you ask for money?

OpenStudy (anonymous):

bahrom, why did you ask for money?

OpenStudy (anonymous):

do you want a real tutoring job? online?

OpenStudy (anonymous):

why didnt you , i mean , ask for money

OpenStudy (anonymous):

hey